OpenStudy (anonymous):

A scale drawing has a scale of 1/4 inch : 12 ft Find the length on the drawing for each actual length. 1. 8ft 2.15ft help!

OpenStudy (radar):

Find out what 1 ft would be: (1/4)/(12)= (1/4)(1/12)=1/48 inch : 1 foot Now 8 ft would be represented by 8 * (1/48) = 1/6 inch on the drawing. 2. 15 times 1/48 = 15/48= 5/16 inch on the drawing.

OpenStudy (anonymous):

i still don't understand ???

OpenStudy (radar):

Do you understand scale? !/4 inch on the drawing represents 12 ft in the reaL world. That is what the scale 1/4 inch : 12 ft means. That was given, that means a line 1/4 inch long represents a dimension of 12 feet. Like on a road map ....undertstand the concept??

OpenStudy (anonymous):

yes i understand that

OpenStudy (radar):

Now carry this a step further, a line 1/2 inch which is twice 1/4 inch would represent a dimension of 24 ft which is twice the 12 ft.

OpenStudy (anonymous):

okay

OpenStudy (radar):

and a line which is 1/8 inch which is 1/2 as long as 1/4 would represent a dimension of 6 feet (1/2 of the 12).

OpenStudy (radar):

You were asked how long is the line on the drawing that would: 1. represent 8 ft and 2. how long would the line be to represent 15 ft.

OpenStudy (anonymous):

yea / sorry i just struggle in math and am frustrated now..

OpenStudy (radar):

Since both dimensions were in feet, it would be easy to solve these to find out how long would a line have to be to represent 1 foot. Then 1. multiply that by 8 solving 1. and multiply by 15 solving 2.

OpenStudy (radar):

The line that would represent 1 foot would be 1/12 as long as the line which represented 12 feet. Well 1/4 in represented 12 feet, so divide 1/4 by 12 getting 1/48 in. which would be equal to 1 foot in the real world. Now you got it made.

OpenStudy (radar):

8 times 1/48= 1/6 for the first 15 times 1/48 =15/48 = 5/16 for the second measurement.

OpenStudy (radar):

Focus and clear the frustration.

OpenStudy (radar):

Don't give up and leave, if you really want help stay with it.

OpenStudy (anonymous):

sry my computer froze up

OpenStudy (radar):

Please review the posts after your previous post.